Thursday, August 2, 2012

Almost there

ODI is now cabled up and cold.  We spent the day watching the vacuum and temperatures carefully as the instrument cooled down.  This is the first time we will leave it cold while we are not around, so Daniel is briefing Krissy, the Observing Assistant who will be on duty tonight, how to do an emergency shutdown if necessary.  The rest of us will depart and return tomorrow for a day of biases and dome flats.  If all goes well, we will be opening the dome next week. 

Daniel and I worked out a schedule today for the engineering verification activities.  Since we are still in the monsoon, we will take advantage of the clear nights when they occur, and otherwise work on the stuff that can be done in the daytime.  All this is subject to change as we proceed, but the priorities are:

Aug 6-10: focus sequence; manual guide star acquisition; focus w/ guide star; telescope guiding; telescope telemetry

Aug 13-17: focus sensors; world coordinate system; on-axis wavefront adjustment; filter offsets

Aug 20-24: Image characterization; science GUI; data for PPA

Aug 27-31: ADC; begin photometric characterization; OT preliminary testing

Sept 3-7: photometric characterization

Sept 10-17: wide-field wavefront adjustment; image quality analysis; update exposure time calculator

It will be a busy period, but this should put us in good shape to begin scientific commissioning in early October.


No comments:

Post a Comment